7) NaI + Br2 à NaBr + I2​​​Mass of NaI: 10.0g​
Experimental Mass of Br2: 4.52g​​Experimental Yield of NaBr:​6.14g ​​Experiment Yield of I2: 6.93g
1
Expert's answer
2016-11-08T14:56:08-0500
8) Molar mass of NaI is 149.9 g/mol

Molar mass of Br2 is 159.8 g/mol

Molar mass of NaBr is 102.89 g/mol

Molar mass of I2 is 253.82 g/mol

n = m/M

n(NaI)=10.0/149.9=0.07 mol

n (Br2) = 4.52/159.8 = 0.03 mol

n (NaBr) = 6.14/102.89 = 0.06 mol

n (I2) = 6.93/253.82 = 0.03 mol

According to this, let’s put coefficients to the equation:

2NaI + Br2  2NaBr + I2